Twice the area of a square is 72 miles. What is the length of each side of the square?
Answer by checkley71(8403)

2X^2=72
X^2=72/2
X^2=36
X=SQRT36
X=6 ANSWER.
PROOF
2*6^2=72
2*36=72
72=72